# Primordial gravitational waves in Wheeler-DeWitt non-commutative linearized branch-cut quantum gravity

### 12 Jan 2024 v1

@@ IX Acknowledgements
$\Biggl{(}\frac{\alpha_{T}\,\kappa_{\infty}}{1+\alpha_{T}}\Biggr{)}^{2}\approx 1.557411\times 10^{-6}\,,$ (155)
@@
$g_{*}=g_{*}^{\rm{[SM]}}=100,\quad\mbox{and thus}\quad\Biggl{(}\frac{100}{g_{*}}\Biggr{)}^{1/3}=1\,.$ (156)
@@
$\Bigl{(}\frac{0.11u^{3}_{\omega}}{0.42+u^{2}_{\omega}}\Bigr{)}\approx 7.746479\times 10^{-2}\,,$ (157)
@@
$\Bigl{(}\frac{0.62}{1.8+u^{2}_{\omega}-0.1n_{\omega}}\Bigr{)}\approx 2.29629\times 10^{-1}\,.$ (158)
@@
$V(T,\phi)=\frac{\gamma}{2}\Bigl{(}T^{2}-T_{0}^{2}\Bigr{)}\phi^{2}-\frac{\sigma}{3}T\phi^{3}+\frac{\lambda}{4}\phi^{4}\,.$ (159)
@@
$\Biggl{(}\frac{\beta}{H_{*}}\Biggr{)}\simeq 4\,\ln\Biggl{(}\frac{M_{P}}{T_{*}}\Biggr{)}\,.$ (160)
@@
Appendix C
In this appendix we briefly describe the steps to obtain the gravitational waves spectra described by the spectral shape of the signal, considering the sub-horizon condition with a matter-energy source given by equation (115). To solve this equation, due to the technical difficulties of its resolution, as we stressed before, we used powerful computational algorithms, based on the Runge-Kutta-Fehlberg method, that made it possible to obtain algebraic solutions without the need to implement simplifications or any computational approximation. The corresponding solution of the field equation that describes the evolution over time of the metric disturbance having as its source the potential defined in equation (93), that configures the dynamic composition of matter in the primordial cosmic period, is
${h}(k,\tau)$ $=$ $c_{2}(\eta)\sin(k\tau)+c_{1}(\eta)\cos(k\tau)+\frac{1}{k^{4}\eta^{3}}\Bigl{[}k^{2}\Bigl{(}6.28319\tau^{2}\eta^{2}+\tau\bigl{(}-4.18879\eta^{5/2}+16.7384\eta^{6}-23.4572\eta^{5}$ (161)
-- $16.7552\eta^{4}-6.70206\eta^{2}-0.150796\bigr{)}+50.2655\eta^{7/2}\Bigr{)}-12.5664\eta^{2}\Bigr{]}\,.$
$S_{b}(\nu/\nu_{c})$, the spectral shape of the signal, is determined in terms of the solution (161) of equation (115), after integrating out the time dependence of $S_{b}(\nu/\nu_{c},\tau)$, taking the limits 00 to $\beta/H_{*}$, with $k\to\nu/\nu_{c}$:
@@
$S_{n}(\nu/\nu_{c})$$\approx$ $\Bigl{(}\sin\Bigl{(}k\frac{\beta}{H_{*}}\Bigr{)}+\cos\Bigl{(}k\frac{\beta}{H_{*}}\Bigr{)}\Bigr{)}\Biggl{(}0.0795775\Bigl{(}\frac{\nu_{c}}{\nu}\frac{H_{*}}{\beta}\Bigr{)}-0.159155\Bigl{(}\frac{\nu_{c}}{\nu}\frac{H_{*}}{\beta}\Bigr{)}^{\!\!3}+1.90986\Bigl{(}\frac{\nu_{c}}{\nu}\frac{H_{*}}{\beta}\Bigr{)}^{\!\!5}\!\Biggr{)}+{\cal O}\Bigl{(}\Bigl{(}\frac{H_{*}}{\beta}\Bigr{)}^{6}\Bigr{)}$ (163)
$+$ $\Bigl{(}\frac{\nu_{c}}{\nu}\Bigr{)}^{5}\Biggl{(}3.15843\Bigl{(}\frac{\beta}{H_{*}}\Bigr{)}+92.4467\Bigl{(}\frac{\beta}{H_{*}}\Bigr{)}^{1/2}+169.119log\Bigl{(}\frac{\beta}{H_{*}}\Bigr{)}+0.150401\Bigl{(}\frac{H_{*}}{\beta}\Bigr{)}+0.733704\Bigl{(}\frac{H_{*}}{\beta}\Bigr{)}^{3/2}$
-- $0.000596831\Bigl{(}\frac{H_{*}}{\beta}\Bigr{)}^{3}\Biggr{)}+0.0795775log\Bigl{(}\frac{\nu_{c}}{\nu}\Bigr{)}-0.0397887log\Bigl{(}\Bigl{(}\frac{\nu_{c}}{\nu}\Bigr{)}^{2}\Bigr{)}+0.125+{\cal O}\Bigl{(}\Bigl{(}\frac{H_{*}}{\beta}\Bigr{)}^{11/2}\Bigr{)}$
$+$ $\Bigl{(}\sin\Bigl{(}k\frac{\beta}{H_{*}}\Bigr{)}-\cos\Bigl{(}k\frac{\beta}{H_{*}}\Bigr{)}\Bigr{)}\Biggl{(}0.0795775\Bigl{(}\frac{\nu_{c}}{\nu}\frac{H_{*}}{\beta}\Bigr{)}^{2}-0.477465\Bigl{(}\frac{\nu_{c}}{\nu}\frac{H_{*}}{\beta}\Bigr{)}^{4}+{\cal O}\Bigl{(}\Bigl{(}\frac{H_{*}}{\beta}\Bigr{)}^{6}\Biggr{)}$
$+$ $\Bigl{(}\sin\Bigl{(}k\frac{\beta}{H_{*}}\Bigr{)}+\cos\Bigl{(}k\frac{\beta}{H_{*}}\Bigr{)}\Bigr{)}\Biggl{(}-0.0795775\Bigl{(}\frac{\nu_{c}}{\nu}\frac{H_{*}}{\beta}\Bigr{)}+{\cal O}\Bigl{(}\Bigl{(}\frac{H_{*}}{\beta}\Bigr{)}^{6}\Biggr{)}\Biggr{)}\,.$

### 03 Feb 2024 v2

@@ Appendix B Generation of gravitational waves: bubble collisions
$\Biggl{(}\frac{\alpha_{T}\,\kappa_{\infty}}{1+\alpha_{T}}\Biggr{)}^{2}\approx 1.557411\times 10^{-6}\,,$ (B.1)
@@
$g_{*}=g_{*}^{\rm{[SM]}}=100,\quad\mbox{and thus}\quad\Biggl{(}\frac{100}{g_{*}}\Biggr{)}^{1/3}=1\,.$ (B.2)
@@
$\Bigl{(}\frac{0.11u^{3}_{\omega}}{0.42+u^{2}_{\omega}}\Bigr{)}\approx 7.746479\times 10^{-2}\,,$ (B.3)
@@
$\Bigl{(}\frac{0.62}{1.8+u^{2}_{\omega}-0.1n_{\omega}}\Bigr{)}\approx 2.29629\times 10^{-1}\,.$ (B.4)
@@
$V(T,\phi)=\frac{\gamma}{2}\Bigl{(}T^{2}-T_{0}^{2}\Bigr{)}\phi^{2}-\frac{\sigma}{3}T\phi^{3}+\frac{\lambda}{4}\phi^{4}\,.$ (B.5)
@@
$\Biggl{(}\frac{\beta}{H_{*}}\Biggr{)}\simeq 4\,\ln\Biggl{(}\frac{M_{P}}{T_{*}}\Biggr{)}\,.$ (B.6)
@@
Appendix C Gravitational waves spectral shape of the signal
In this appendix we briefly describe the steps to obtain the gravitational waves spectra described by the spectral shape of the signal, considering the sub-horizon condition with a matter-energy source given by equation (6.1). To solve this equation, due to the technical difficulties of its resolution, as we stressed before, we used powerful computational algorithms, based on the Runge-Kutta-Fehlberg method, that made it possible to obtain algebraic solutions without the need to implement simplifications or any computational approximation. The corresponding solution of the field equation that describes the evolution over time of the metric disturbance having as its source the potential defined in equation (5.20), that configures the dynamic composition of matter in the primordial cosmic period, is
[…]
$-23.4572\eta^{5}-16.7552\eta^{4}-6.70206\eta^{2}-0.150796\bigr{)}+50.2655\eta^{7/2}\Bigr{)}-12.5664\eta^{2}\Bigr{]}.$
$S_{b}(\nu/\nu_{c})$, the spectral shape of the signal, is determined in terms of the solution (LABEL:equationh*) of equation (6.1), after integrating out the time dependence of $S_{b}(\nu/\nu_{c},\tau)$, taking the limits 00 to $\beta/H_{*}$, with $k\to\nu/\nu_{c}$:
@@ Appendix C Gravitational waves spectral shape of the signal
$S_{n}(\nu/\nu_{c})$$\approx$ $\Bigl{(}\sin\Bigl{(}k\frac{\beta}{H_{*}}\Bigr{)}+\cos\Bigl{(}k\frac{\beta}{H_{*}}\Bigr{)}\Bigr{)}\Biggl{\{}0.0795775\Bigl{(}\frac{\nu_{c}}{\nu}\frac{H_{*}}{\beta}\Bigr{)}-0.159155\Bigl{(}\frac{\nu_{c}}{\nu}\frac{H_{*}}{\beta}\Bigr{)}^{\!\!3}$
$+1.90986\Bigl{(}\frac{\nu_{c}}{\nu}\frac{H_{*}}{\beta}\Bigr{)}^{\!\!5}\!\Biggr{)}+{\cal O}\Bigl{(}\Bigl{(}\frac{H_{*}}{\beta}\Bigr{)}^{6}\Bigr{)}+\Bigl{(}\frac{\nu_{c}}{\nu}\Bigr{)}^{5}\Biggl{[}3.15843\Bigl{(}\frac{\beta}{H_{*}}\Bigr{)}+92.4467\Bigl{(}\frac{\beta}{H_{*}}\Bigr{)}^{1/2}$
$+169.119log\Bigl{(}\frac{\beta}{H_{*}}\Bigr{)}+0.150401\Bigl{(}\frac{H_{*}}{\beta}\Bigr{)}+0.733704\Bigl{(}\frac{H_{*}}{\beta}\Bigr{)}^{3/2}\!-0.000596831\Bigl{(}\frac{H_{*}}{\beta}\Bigr{)}^{3}\!\Biggr{]}$
$+0.0795775log\Bigl{(}\frac{\nu_{c}}{\nu}\Bigr{)}-0.0397887log\Bigl{(}\Bigl{(}\frac{\nu_{c}}{\nu}\Bigr{)}^{2}\Bigr{)}+0.125+{\cal O}\Bigl{(}\Bigl{(}\frac{H_{*}}{\beta}\Bigr{)}^{11/2}\Bigr{)}$
$+\Biggl{[}\Bigl{(}\sin\Bigl{(}k\frac{\beta}{H_{*}}\Bigr{)}-\cos\Bigl{(}k\frac{\beta}{H_{*}}\Bigr{)}\Bigr{)}\Biggl{(}0.0795775\Bigl{(}\frac{\nu_{c}}{\nu}\frac{H_{*}}{\beta}\Bigr{)}^{2}-0.477465\Bigl{(}\frac{\nu_{c}}{\nu}\frac{H_{*}}{\beta}\Bigr{)}^{4}\Biggr{)}$
$+\Bigl{(}\sin\Bigl{(}k\frac{\beta}{H_{*}}\Bigr{)}+\cos\Bigl{(}k\frac{\beta}{H_{*}}\Bigr{)}\Bigr{)}\Biggl{(}-0.0795775\Bigl{(}\frac{\nu_{c}}{\nu}\frac{H_{*}}{\beta}\Bigr{)}\Biggr{)}+{\cal O}\Bigl{(}\Bigl{(}\frac{H_{*}}{\beta}\Bigr{)}^{6}\Biggr{]}\Biggr{\}}.$